OpenStudy (anonymous):

Solve the equation below for x: 3(x+1)=8(x-1)+5 A. x = -3/5 B. x = 6/5 C. x = 5/6 D. x = -5/4

OpenStudy (anonymous):

B

OpenStudy (anonymous):

3x + 3 = 8x-8 + 5 3x+3 = 8x-3 3x+6 = 8x 6 = 5x x = 6/5 B